OpenStudy (anonymous):

MEDAL AND FAN Solve 5^(2x) = 7^(x−1)

OpenStudy (anonymous):

answer choices 1.52864 -1 -1.52864 1

OpenStudy (anonymous):

You have to take Logarithm on both sides...

OpenStudy (anonymous):

Take Log to base 10 to both sides..

OpenStudy (anonymous):

so the answer is 1

OpenStudy (anonymous):

Is it? Show me how you got?

OpenStudy (anonymous):

Log 10 5^(2x)= LOG 10 7^(x-1)

OpenStudy (anonymous):

Yes, what you did next?

OpenStudy (anonymous):

Use power rule of Logarithms which is: \[\log(m^n) = n \cdot \log(m)\]

OpenStudy (anonymous):

so then the answer 1.52864

OpenStudy (anonymous):

Really? You can show all the steps you did? or wait, let me check that once..

OpenStudy (anonymous):

i skip some step but i feel like that the answer

OpenStudy (anonymous):

You feel or you have solved it?

OpenStudy (anonymous):

i feel like it is the answer

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

A house cost $120,000 when it was purchased. The value of the house increases by 10% each year. Find the rate of growth each month and select the correct answer below.

OpenStudy (anonymous):

0.797% 0.00797% 0.833% 0.00833%

OpenStudy (anonymous):

is the answer the first one?